A body of mass 2.60 kg is pushed straight upward by a 31.2 N vertical force. What is it’s acceleration?
1
Expert's answer
2019-09-24T12:11:13-0400

Newton's second Law is applied to the body.


"F_{Net}=m*a"


Where

  • The object's mass "m=2.60kg"
  • Acceleration a"a"


But the net force is equal to the sum of the forces


"F_{Net}=\sum F \\ F_{net}=F-W"


Where:

External force "F=31.2Nw"

The weight is "W=m*g \\ W=2.60kg*9.81\frac{m}{s^{2}}\\W=25.5Nw"


Combining expressions


"m*a=F-W"


Expression for acceleration.

"m*a=F-W \\a=\frac{F-W}{m}"


Numerically evaluating


"a=\frac{32.5Nw-25.5Nw}{2.60kg}=2.69\frac{m}{s^{2}}"


The acceleration of the object is

"\boxed{a=2.69\frac{m}{s^{2}}}"